# A simple controller design to achieve iso-damping robustness: Non-iterative data-driven approach based on fractional-order reference model

### v1

@@ Abstract
thanks: This paper was not presented at any IFAC meeting. This work was supported in part by JSPS KAKENHI Grant Number JP23K19084. Corresponding author A. Yonezawa.
1 Introduction
1.1 Motivation
@@
1.2 Related works
Controller synthesis based on a fractional-order calculus (FC) has generated a great deal of attention: fractional-order (FO) system and control [23][41][45][57]. FC is a mathematical framework generalizing traditional integer-order (IO) calculus, which considers non-integer order differentiations and integrations. Owing to its design flexibility, FO control is a powerful and effective approach to achieving good control performance and robustness. In particular, the Bode’s ideal transfer function (BITF) is an FO transfer function with the ideal flat-phase specification; it is often utilized for controller design to achieve the iso-damping robustness [4].
Various approaches have been proposed to achieve iso-damping on the basis of FO control. These methods aim to tune some design parameters of the fixed structure controller, such as the proportional-integral-derivative (PID) controller and the FO-PID controller, so as to realize the flat-phase property. The analytical conditions for the flat-phase have been derived for a special class of plants and controllers: the FO-PID controller for simplified FO plants with time delay [58], the FO-PID controller for stable first-order plus time delay (FOPTD) plant [9], and the FO-PID controller with FO lowpass filter [10]. The graphical approach (e.g., plotting the feasible parameter region or achievable performance specification) is an effective way to determine controller parameters satisfying the analytical flat-phase condition. As for the graphical approach, one study has utilized the Bode phase plot [29], one has proposed a flat-phase curve [11], and a novel procedure to tune the FO-[PID] controller has been established for the FOPTD plant [47]. It should be noted that both the analytical and graphical approaches require accurate mathematical models and parameters of the actual controlled plant. However, the necessity of the plant model and parameters imposes heavy burdens on designers and causes the technique to be fail under modeling errors. Moreover, in most cases, the analytical and graphical policies are derived for very simple plants such as the FOPTD system. Thus, not only the plant model but also its bold approximation is necessary to employ such a method. Controller design based on the approximated model often causes severe performance degradation due to the gap between the actual plant and the reduced one for controller design.
@@ 1.2 Related works
Recently, controller design techniques based on the data rather than the mathematical model of the plant have attracted a great deal of attention: direct data-driven control (D3C) [6][25][44][46]. Exploiting the D3C approach to design an FO control system can overcome the drawbacks of the abovementioned challenges, since these problems are mainly raised by the needs for accurate plant models and parameters. Several studies have employed frequency response data for designing FO controllers achieving the iso-damping property [38][39][49]. However, obtaining the frequency response is cumbersome and often requires a special experimental setup.
Model-reference (MR) D3C [33] is a simple and practical framework for achieving the desired closed-loop characteristics while avoiding burdensome frequency response computation. Various MR-D3C techniques have been examined for IO control systems: iterative feedback tuning (IFT) [19], virtual reference feedback tuning (VRFT) [8][14][18], correlation-based tuning [43], fictitious reference iterative tuning (FRIT) [22], to name a few. Many studies have reported the effectiveness of MR-D3C in the IO control framework [32][51][52]. For example, the VRFT approach has been employed for the vehicle yaw rate control [53][55].
On the other hand, to the best of our knowledge, few studies have explored an MR-D3C approach for FO control. The IFT technique has been employed to tune the FO-PID controller for a hydraulic actuator [24]. However, the tuning scheme presented in [24] requires multiple experiments, imposing heavy burdens on designers. The FO-PID controller has been tuned on the basis of the VRFT [1][2][48]. Nevertheless, the conventional VRFT does not consider closed-loop stability. Although one study has tuned the FO-PID controller based on the MR-D3C approach, it does not consider the iso-damping robustness [56]. Consequently , to the best of our knowledge, no studies have proposed simple and practical FO MR-D3C techniques especially for achieving the iso-damping property.
1.3 Contribution and novelty
@@
    (Simplicity) The proposed approach does not require the mathematical model of the actual controlled plant, providing freedom from costly and burdensome modeling procedures. Therefore, the present approach is simpler than model-based approaches such as analytical and graphical approaches [9][10][11][23][29][47][58] and numerical optimization based on the plant model [4][21][26].
@@
    (Reliability) Unlike the conventional FO MR-D3C techniques [1][2][48], information regarding the closed-loop pole is explicitly reflected in the proposed optimization problem for the controller design. This feature allows the present technique to explicitly consider the bounded-input bounded-output (BIBO) stability of the closed-loop system from a bounded reference input to a bounded output. Therefore, the proposed design approach yields a reliable controller.
@@
Owing to contributions (1)–(5), we can easily incorporate useful and important robustness, i.e., the iso-damping, into various control systems. This study contributes to improve the safety, reliability, and performance of a wide variety of control systems in practice.
1.4 Structure of this paper
This paper is organized as follows. Section 2 summarizes some preliminaries and states the problem addressed in this study. Section 3 describes the proposed controller design approach. We show the numerical studies to demonstrate the validity of the proposed approach in Section 4. The discussion is given in Section 5. Section 6 concludes the paper.
2 Preliminaries
@@
2.1 The flat-phase property and iso-damping
@@
2.2 BIBO stability of proper rational discrete-time IO SISO LTI system
@@
2.3 Overview of the controller design problem in this study
@@
3 The proposed approach
This section describes the controller tuning problem described in Section 2.3 so as to achieve the iso-damping robust control.
3.1 Robust controller design via MR-D3C framework
@@
Then, the controller design problem in Section 2.3 can be formally formulated as follows:
@@
Since $L_{flat}\left(z\right)$ is constructed based on the FO transfer function, it is preferable that the class of the controller is also chosen as the FC-based controller such as $C\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FOPID}\left(s;\theta\right)\right)$, where $C_{FOPID}\left(s;\theta\right)$ is the standard FO-PID controller with the tunable parameter $\theta$. This fact is demonstrated in Example 1 in Section 4. $\blacksquare$
@@
3.2 Fictitious reference signal
The proposed approach addresses Problem 1 based on the fictitious reference signal [3][22][36]. The fictitious reference signal $\tilde{r}_{[0:N]}\left(\theta\right)=\left\{\tilde{r}_{k}\left(\theta\right)\right\}_{k=0}^{N}$ is defined as
@@ 3.2 Fictitious reference signal
One study has proposed a MR-D3C technique based on the fictitious reference signal: FRIT [22]. However, the controller provided by FRIT may destabilize the closed-loop system as pointed out in [37][50]. To overcome this difficulty, several studies have examined to modify FRIT for IO control [37][50]. $\blacksquare$
3.3 The proposed controller design approach
@@
Let $M_{ref}\left(z\right)$ be the reference model constructed based on the FO open-loop transfer function with the flat-phase property, as stated in Section 3.1. Let $m^{ref}=\left\{m_{k}^{ref}\right\}_{k=0}^{\infty}$ be the impulse response sequence of $M_{ref}\left(z\right)$ and $m_{[0:N]}^{ref}\coloneqq\left\{m_{k}^{ref}\right\}_{k=0}^{N}$. Then, the solution of Problem 1 is given by
@@ 3.4 Stability analysis
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI transfer function with the minimal state-space realization $(A,B,C,D)$, i.e., $G(z)=C\left(zI-A\right)^{-1}B+D$ ($I$ is the identity matrix with the compatible dimension). The impulse response sequence of $G\left(z\right)$ is denoted as $g=\left\{g_{k}\right\}_{k=0}^{\infty}$, where $g_{0}=D$ and $g_{k}=CA^{k-1}B$ for $k\geq 1$. Then,
  1. (F1)
    $G\left(z\right)$ is BIBO stable if and only if $\rho\left(A\right)<1$;
  2. (F2)
    $G\left(z\right)$ is BIBO stable if and only if $\left(\left\lVert g\right\rVert_{2}\right)<\infty$. $\blacksquare$
(Proof) (F1) Note that $\Lambda\left(A\right)$ coincides the poles of $G\left(z\right)$ including multiplicity due to the minimality of $(A,B,C,D)$ (Theorem 12.9.16 in [5]). Then, the conclusion follows from Proposition 1. (F2) Assume that $G\left(z\right)$ is BIBO stable. Then, due to (C1) in Proposition 1, $\left(\left\lVert g\right\rVert_{2}\right)^{2}<\left\lVert g\right\rVert_{\infty}\left\lVert g\right\rVert_{1}<\infty$ showing the sufficiency. In contrast, suppose that $G\left(z\right)$ is not BIBO stable, implying $\rho\left(A\right)\geq 1$. Here, for $k\to\infty$, $\left\lvert g_{k}\right\rvert^{2}=\left\lvert CA^{k-1}B\right\rvert^{2}\to 0$ if and only if $A^{k-1}\to 0$. However, $A^{k-1}\to 0$ if and only if $\rho\left(A\right)<1$ (Theorem 5.6.12 in [20]), proving that $\left\lVert g\right\rVert_{2}$ does not converge to finite value. ∎
@@
$\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)\\ .$ (12)
That is, Here, $J\left(\theta^{\ast}\right)=\left(\left\lVert\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right\rVert_{2}\right)^{2}=\left(\left\lVert\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}\right)^{2}$ due to (8) and (12). Therefore, where $\beta_{RD}\in\mathbb{R}_{+}$ denotes the induced 2-norm of $\left(R^{D}\right)^{-1}$. Here, there exists a finite $\beta_{ref}\in\mathbb{R}_{+}$ such that $\lim_{N\to\infty}\left\lVert\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}=\left\lVert m^{ref}\right\rVert_{2}<\beta_{ref}$ due to Assumption 1 and Lemma 2. Finally, which proves the conclusion. ∎
@@
3.5 Summary of the proposed controller design
@@
In practice, the IO approximation and discretization have to be carried out so as to implement an FO controller [13][41]. The IO approximation and discretization may affect the closed-loop performance and stability [16]. Unlike conventional approaches, the proposed design scheme explicitly handles the IO-approximated and discretized controller. That is, the proposed approach evaluates the resultant closed-loop performance and stability for the IO-approximated and discretized, i.e., ready-to-implement, controller, showing the reliability of the proposed approach. $\blacksquare$
4 Numerical examples
@@
4.1 Example 1: High-order process model (FO control vs. IO control)
@@
where $K_{p},K_{i},K_{d},K_{fp},K_{fi},K_{fd},\lambda,\mu,\tau_{i},\tau_{f}\in\mathbb{R}_{+}\cup\left\{0\right\}$. In this study, we set $\tau_{i}=\tau_{f}=t_{s}$. The tuning conditions are listed in Table 1. The IO approximation is performed via the Oustaloup recursive filter (order: 5; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-4},10^{4}\right)\;\mathrm{rad/s}$) for both $C_{FOPID}\left(z;\theta\right)$ and $M_{ref}\left(z\right)$. In the initial data correction, the simulation time is $40\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data for the controller tuning are the same to each other in the FO-PID controller and the IO-PID controller, since $C_{IOPID}\left(z;\theta_{0}\right)=C_{FOPID}\left(z;\theta_{0}\right)$. Fig. 4 shows the input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ for tuning both $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$.
Table 2 summarizes the tuning results for $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$ provided by the proposed approach (iso-IDFRIT). Table 3 lists the phase margin $\phi_{m}$ and the gain crossover frequency $\omega_{c}$ provided by the tuned controllers. Fig. 5 demonstrates the control results obtained by the resultant controllers. The Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ are shown in Fig. 6. In Fig. 5, the black line indicates the setpoint reference input. In Figs. 5 and 6, the green, blue, and red lines demonstrate the results due to the reference model ($M_{ref}\left(z\right)$ in Fig. 5 and $L_{flat}\left(z\right)$ in Fig. 6), $C_{IOPID}\left(z;\theta^{\ast}\right)$, and $C_{FOPID}\left(z;\theta^{\ast}\right)$, respectively. Tables 2 and 3 and Figs. 5 and 6 clearly show that the proposed approach successfully provides the IO- and FO-PID controllers yielding the desired closed-loop property. In particular, the closed-loop system due to the FO-PID controller achieves the closer specifications ($\phi_{m}$, $\omega_{c}$, the flat-phase) to the desired ones than that due to the IO-PID controller.
@@
Table 1: Conditions for Example 1
$\theta_{0}$ IO $\begin{bmatrix}K_{p}&K_{i}&K_{d}\\ \end{bmatrix}=\begin{bmatrix}1&0&0\\ \end{bmatrix}$
FO $\begin{bmatrix}K_{fp}&K_{fi}&\lambda&K_{fd}&\mu\\ \end{bmatrix}=\begin{bmatrix}1&0&1&0&1\\ \end{bmatrix}$
Search range IO $K_{p},K_{i},K_{d}\in\begin{bmatrix}0&5\\ \end{bmatrix}$
FO $K_{fp},K_{fi},K_{fd}\in\begin{bmatrix}0&5\\ \end{bmatrix}$, $\lambda,\mu\in\begin{bmatrix}0&2\\ \end{bmatrix}$
@@
Table 2: Tuning results (Example 1)
@@
Table 3: Frequency domain property provided by the tuned controllers (Example 1).
@@
Figure 4: Initial input/output data for the proposed controller tuning (Example 1)
@@
Figure 6: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ (Example 1): (a) overview, and (b) enlarged phase plots around 1 Hz.
@@
4.2 Example 2: Superiority over the conventional approach based on the reduced model
One representative virtue of the proposed approach is the freedom from the necessity of the plant model and the model reduction. This feature allows the proposed approach to achieve better control performance and robustness than the reduced-model-based controller design scheme. Note that almost all conventional analytical and graphical design methodologies rely on the reduced order model, as discussed in Section 1.2. We demonstrate the superiority and robustness of the proposed approach, compared to the traditional iso-damping controller design based on the reduced model. Throughout this example, the Oustaloup recursive filter is employed for the IO approximation (order: 7; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-3},10^{6}\right)\;\mathrm{rad/s}$) for both controllers and reference models. We discretize the plant, the reference model, and the controller in the sampling time $t_{s}=10^{-2}\,s$.

### v2

# Simple Controller Design to Achieve Iso-Damping Robustness: Non-Iterative Data-Driven Approach Based on Fractional-Order Reference Model

@@ Abstract
I Introduction
I-A Motivation
@@
I-B Related works
Control systems involving a fractional-order calculus (FC) has generated a great deal of attention: fractional-order (FO) system theory (e.g., stabilization of FO systems [5, 6, 7, 8, 9], stability tests [10, 11, 12, 13], the generalized KYP lemma for singular FO systems [14], the FC-based gradient algorithm [15]). FC is a mathematical framework generalizing traditional integer-order (IO) calculus, which considers non-integer order differentiations and integrations. Focusing on its rich expression ability and generality, Magin et al. have discussed the benefit of exploiting the idea of FC in the cybernetics field [16]. In particular, FO control is a powerful and effective approach to achieving good control performance and robustness [17]. For example, the Bode’s ideal transfer function (BITF) is an FO transfer function with the ideal flat-phase specification; it is often utilized for controller design to achieve the iso-damping robustness [18]. FC-based robust control is one important direction of the exploitation of FC in the cybernetics field.
Various approaches have been proposed to achieve iso-damping on the basis of FO control. These methods aim to tune some design parameters of the fixed structure controller, such as the FO proportional-integral-derivative (PID) controller, so as to realize the flat-phase property. The analytical conditions for the flat-phase have been derived for a special class of plants and controllers: the FO-PID controller for simplified FO plants with time delay [19], the FO-PID controller for stable first-order plus time delay (FOPTD) plant [20], and the FO-PID controller with FO lowpass filter [21]. The graphical approach (e.g., plotting the feasible parameter region or achievable performance specification) is an effective way to determine controller parameters satisfying the analytical flat-phase condition. As for the graphical approach, one study has utilized the Bode phase plot [22], one has proposed a flat-phase curve [23], and a novel procedure to tune the FO-[PID] controller has been established for the FOPTD plant [24]. It should be noted that both the analytical and graphical approaches require accurate mathematical models and parameters of the actual controlled plant. However, the necessity of the plant model and parameters imposes heavy burdens on designers and causes the technique to be fail under modeling errors. Moreover, in most cases, the analytical and graphical policies are derived for very simple plants such as the FOPTD system. Thus, not only the plant model but also its bold approximation is necessary to employ such a method. Controller design based on the approximated model often causes severe performance degradation due to the gap between the actual plant and the reduced one for controller design.
@@ I-B Related works
Recently, controller design techniques based on the data rather than the mathematical model of the plant have attracted a great deal of attention: direct data-driven control (D3C) [28, 29, 30]. Exploiting the D3C approach to design an FO control system can overcome the drawbacks of the abovementioned challenges, since these problems are mainly raised by the needs for accurate plant models and parameters. Several studies have employed frequency response data for designing FO controllers achieving the iso-damping property [31, 32]. However, obtaining the frequency response is cumbersome and often requires a special experimental setup.
Model-reference (MR) D3C is a simple and practical framework for achieving the desired closed-loop characteristics while avoiding burdensome frequency response computation. Various MR-D3C techniques have been examined for IO control systems: iterative feedback tuning (IFT) [33], virtual reference feedback tuning (VRFT) [34, 35], correlation-based tuning [36], fictitious reference iterative tuning (FRIT) [37], to name a few. Many studies have reported the effectiveness of MR-D3C in the IO control framework [38, 39, 40, 41]. For example, the VRFT approach has been employed for the vehicle yaw rate control [42, 43].
On the other hand, to the best of our knowledge, few studies have explored an MR-D3C approach for FO control. The IFT technique has been employed to tune the FO-PID controller for a hydraulic actuator [44]. However, the tuning scheme presented in [44] requires multiple experiments, imposing heavy burdens on designers. The FO-PID controller has been tuned on the basis of the VRFT [45, 46, 47]. Nevertheless, the conventional VRFT does not consider closed-loop stability. Although one study has tuned the FO-PID controller based on the MR-D3C approach, it does not consider the iso-damping robustness [48]. Consequently , to the best of our knowledge, no studies have proposed simple and practical FO MR-D3C techniques especially for achieving the iso-damping property.
I-C Contribution and novelty
@@
    (Simplicity) The proposed approach does not require the mathematical model of the actual controlled plant, providing freedom from costly and burdensome modeling procedures. Therefore, the present approach is simpler than model-based approaches such as analytical and graphical approaches [19, 20, 21, 22, 23, 24] and numerical optimization based on the plant model [18, 27, 25].
@@
    (Reliability) Unlike the conventional FO MR-D3C techniques [45, 46, 47], information regarding the closed-loop pole is explicitly reflected in the proposed optimization problem for the controller design. This feature allows the present technique to explicitly consider the bounded-input bounded-output (BIBO) stability of the closed-loop system from a bounded reference input to a bounded output. Therefore, the proposed design approach yields a reliable controller.
@@
Owing to contributions (1)–(5), we can easily incorporate useful and important robustness, i.e., the iso-damping, into various control systems. This study contributes to improve the safety, reliability, and performance of a wide variety of control systems in practice. Moreover, this study has the importance from the viewpoint of the incorporation of FC into the cybernetics field. Specifically, this study is a pioneering work of the exploration the interdisciplinary area between data-driven control and the FO system theory.
I-D Structure of this paper
This paper is organized as follows. Section II summarizes some preliminaries and states the problem addressed in this study. Section III describes the proposed controller design approach. We show the numerical studies to demonstrate the validity of the proposed approach in Section IV. The discussion is given in Section V. Section VI concludes the paper.
II Preliminaries
@@
II-A The flat-phase property and iso-damping
@@
II-B BIBO stability of proper rational discrete-time IO SISO LTI system
@@
II-C Overview of the controller design problem in this study
@@
III The proposed approach
This section describes the controller tuning problem described in Section II-C so as to achieve the iso-damping robust control.
III-A Robust controller design via MR-D3C framework
@@
Then, the controller design problem in Section II-C can be formally formulated as follows:
@@
Since $L_{flat}\left(z\right)$ is constructed based on the FO transfer function, it is preferable that the class of the controller is also chosen as the FC-based controller such as $C\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FOPID}\left(s;\theta\right)\right)$, where $C_{FOPID}\left(s;\theta\right)$ is the standard FO-PID controller with the tunable parameter $\theta$. This fact is demonstrated in Example 1 in Section IV. $\blacksquare$
@@
III-B Fictitious reference signal
The proposed approach addresses Problem 1 based on the fictitious reference signal [37, 51]. The fictitious reference signal $\tilde{r}_{[0:N]}\left(\theta\right)=\left\{\tilde{r}_{k}\left(\theta\right)\right\}_{k=0}^{N}$ is defined as
@@ III-B Fictitious reference signal
One study has proposed a MR-D3C technique based on the fictitious reference signal: FRIT [37]. However, the controller provided by FRIT may destabilize the closed-loop system as pointed out in [50, 52]. To overcome this difficulty, several studies have examined to modify FRIT for IO control [50, 52]. $\blacksquare$
III-C The proposed controller design approach
@@
Let $M_{ref}\left(z\right)$ be the reference model constructed based on the FO open-loop transfer function with the flat-phase property, as stated in Section III-A. Let $m^{ref}=\left\{m_{k}^{ref}\right\}_{k=0}^{\infty}$ be the impulse response sequence of $M_{ref}\left(z\right)$ and $m_{[0:N]}^{ref}\coloneqq\left\{m_{k}^{ref}\right\}_{k=0}^{N}$. Then, the solution of Problem 1 is given by
@@ III-D Stability analysis
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI transfer function. The impulse response sequence of $G\left(z\right)$ is denoted as $g=\left\{g_{k}\right\}_{k=0}^{\infty}$. Then, $G\left(z\right)$ is BIBO stable if and only if $\left\lVert g\right\rVert_{2}<\infty$. $\blacksquare$
(Proof) Let $(A,B,C,D)$ be the the minimal state-space realization of $G\left(z\right)$, i.e., $G(z)=C\left(zI-A\right)^{-1}B+D$ ($I$ is the identity matrix with the compatible dimension). Then, $g_{0}=D$ and $g_{k}=CA^{k-1}B$ for $k\geq 1$. Note that $\Lambda\left(A\right)$ coincides the poles of $G\left(z\right)$ including multiplicity due to the minimality of $(A,B,C,D)$ (Theorem 12.9.16 in [53]). This implies that $G\left(z\right)$ is BIBO stable if and only if $\rho\left(A\right)<1$ due to (C2) in Proposition 1.
Assume that $G\left(z\right)$ is BIBO stable. Then, due to (C1) in Proposition 1, $\left(\left\lVert g\right\rVert_{2}\right)^{2}<\left\lVert g\right\rVert_{\infty}\left\lVert g\right\rVert_{1}<\infty$. In contrast, suppose that $G\left(z\right)$ is not BIBO stable, i.e., $\rho\left(A\right)\geq 1$. Here, for $k\to\infty$, $\left\lvert g_{k}\right\rvert^{2}=\left\lvert CA^{k-1}B\right\rvert^{2}\to 0$ if and only if $A^{k-1}\to 0$. However, $A^{k-1}\to 0$ if and only if $\rho\left(A\right)<1$ (Theorem 5.6.12 in [54]), proving that $\left\lVert g\right\rVert_{2}$ is not finite. ∎
@@
. (12)
That is, Here, $J\left(\theta^{\ast}\right)=\left(\left\lVert\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right\rVert_{2}\right)^{2}=\left(\left\lVert\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}\right)^{2}$ due to (8) and (LABEL:Eq_Vec-epsilon). Therefore,
$\begin{split}&\left\lVert\mathrm{vec}\left(t_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}\\ &=\left\lVert\left(R^{D}\right)^{-1}\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)+\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}\\ &\leq\beta_{RD}\sqrt{J\left(\theta^{\ast}\right)}+\left\lVert\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}\end{split}$ (14)
where $\beta_{RD}\in\mathbb{R}_{+}$ denotes the induced 2-norm of $\left(R^{D}\right)^{-1}$. Here, there exists a finite $\beta_{ref}\in\mathbb{R}_{+}$ such that $\lim_{N\to\infty}\left\lVert\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}=\left\lVert m^{ref}\right\rVert_{2}<\beta_{ref}$ due to Assumption 1 and Lemma 2. Finally,
$\begin{split}\lim_{N\to\infty}\left\lVert\mathrm{vec}\left(t_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}&=\left\lVert t\left(\theta^{\ast}\right)\right\rVert_{2}\\ &\leq\sqrt{\alpha}\beta_{RD}+\beta_{ref}<\infty\end{split}$ (15)
which proves the conclusion. ∎
@@
III-E Summary of the proposed controller design
@@
In practice, the IO approximation and discretization have to be carried out so as to implement an FO controller [17, 55]. The IO approximation and discretization may affect the closed-loop performance and stability [56]. Unlike conventional approaches, the proposed design scheme explicitly handles the IO-approximated and discretized controller. That is, the proposed approach evaluates the resultant closed-loop performance and stability for the IO-approximated and discretized, i.e., ready-to-implement, controller, showing the reliability of the proposed approach. $\blacksquare$
IV Numerical examples
@@
IV-A Example 1: High-order process (FO control vs. IO control)
@@
where $K_{p},K_{i},K_{d},K_{fp},K_{fi},K_{fd},\lambda,\mu,\tau_{i},\tau_{f}\in\mathbb{R}_{+}\cup\left\{0\right\}$. The FOPID controller is a typical example of the FC-based control scheme [61]. In this study, we set $\tau_{i}=\tau_{f}=t_{s}$. The tuning conditions are listed in Table I. The IO approximation is performed via the Oustaloup recursive filter (order: 5; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-4},10^{4}\right)\;\mathrm{rad/s}$) for both $C_{FOPID}\left(z;\theta\right)$ and $M_{ref}\left(z\right)$. In the initial data correction, the simulation time is $40\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data for the controller tuning are the same to each other in the FO-PID controller and the IO-PID controller, since $C_{IOPID}\left(z;\theta_{0}\right)=C_{FOPID}\left(z;\theta_{0}\right)$. Fig. 4 shows the input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ for tuning both $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$.
Table II summarizes the tuning results for $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$ provided by the proposed approach (iso-IDFRIT). Table III lists the phase margin $\phi_{m}$ and the gain crossover frequency $\omega_{c}$ provided by the tuned controllers. Fig. 5 demonstrates the control results obtained by the resultant controllers. The Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ are shown in Fig. 6. In Fig. 5, the black line indicates the setpoint reference input. In Figs. 5 and 6, the green, blue, and red lines demonstrate the results due to the reference model ($M_{ref}\left(z\right)$ in Fig. 5 and $L_{flat}\left(z\right)$ in Fig. 6), $C_{IOPID}\left(z;\theta^{\ast}\right)$, and $C_{FOPID}\left(z;\theta^{\ast}\right)$, respectively. Tables II and III and Figs. 5 and 6 clearly show that the proposed approach successfully provides the IO- and FO-PID controllers yielding the desired closed-loop property. In particular, the closed-loop system due to the FO-PID controller achieves the closer specifications ($\phi_{m}$, $\omega_{c}$, the flat-phase) to the desired ones than that due to the IO-PID controller.
@@
TABLE I: Conditions for Example 1
$\theta_{0}$ $\begin{bmatrix}K_{p}&K_{i}&K_{d}\\ \end{bmatrix}=\begin{bmatrix}1&0&0\\ \end{bmatrix}$
$\begin{bmatrix}K_{fp}&K_{fi}&\lambda&K_{fd}&\mu\\ \end{bmatrix}=\begin{bmatrix}1&0&1&0&1\\ \end{bmatrix}$
Search range $K_{p},K_{i},K_{d}\in\begin{bmatrix}0&5\\ \end{bmatrix}$
$K_{fp},K_{fi},K_{fd}\in\begin{bmatrix}0&5\\ \end{bmatrix}$,   $\lambda,\mu\in\begin{bmatrix}0&2\\ \end{bmatrix}$
@@
TABLE II: Tuning results (Example 1)
@@
TABLE III: Frequency domain property provided by the tuned controllers (Example 1).
@@
Figure 4: Initial input/output data for the proposed approach (Example 1)
@@
Figure 6: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ (Example 1): (a) overview, and (b) enlarged phase plots around $1\,\mathrm{rad/s}$.
@@
IV-B Example 2: Soft robot (superiority over the conventional model-based approach)
One representative virtue of the proposed approach is the freedom from the necessity of the plant model and the model reduction. This feature allows the proposed approach to achieve better control performance and robustness than the reduced-model-based controller design scheme. Note that almost all conventional analytical and graphical design methodologies rely on the reduced order model, as discussed in Section I-B. We demonstrate the superiority and robustness of the proposed approach, compared to the traditional iso-damping controller design based on the reduced model. Throughout this example, the Oustaloup recursive filter is employed for the IO approximation (order: 7; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-3},10^{6}\right)\;\mathrm{rad/s}$) for both controllers and reference models. We discretize the plant, the reference model, and the controller in the sampling time $t_{s}=10^{-2}\,s$.